Translate the following into a quadratic equation, and solve it: The length of a rectangular garden is four times its width; if the area of the garden is 196 square meters, what are its dimensions
Found 2 solutions by user_dude2008, RAY100: Answer by user_dude2008(1862) 
 "The length of a rectangular garden is four times its width; if the area of the garden is 196 square meters" ----> 196=4x^2
196=4x^2
49=x^2
x=7
Ans: width=7, length=28
Answer by RAY100(1637) 
 Area=Length * Width
but L=4w
a=4w*w=4w^2
given A=196
196=4w^2
49=w^2
w= +/- 7,, but (-7) in not realistic
,
w = 7
and l=4w=28
,
checking
a=lxw=28*7=196,,,ok
| |
|